OpenStudy (anonymous):

How do I find the integral of (1-2x^2)^3(x) dx

OpenStudy (phi):

if it is \[ \int (1-2x^2)^3 x \ dx \] let \[ u = (1-2x^2) \\ du = -4x\ dx\] and \[ x\ dx= -\frac{1}{4} du \] replace \( (1-2x^2) \) with \( u \) replace \(x\ dx \) with \(-\frac{1}{4} du \)

OpenStudy (anonymous):

Thankyou(: